OpenStudy (anonymous):

solve the system of equations: x+2y + z=1 2x+3y+2z=0 -x-3y+3z=1 *****Please Help!!!*****

jhonyy9 (jhonyy9):

so like a first step assum first and 3rd equation so than will be eliminated the terms of x hope you can continue it

OpenStudy (wolf1728):

A) x+2y + z=1 B) 2x+3y+2z=0 C) -x-3y+3z=1 Adding equations A & C -y +4z = 2 Multiplying equation by -1 y -4z = -2 y = (4z -2) Inputting this in equation A x + 2*(4z-2) + z =1 x + 8z -4 + z = 1 x + 9z = 5 multiply by -2 D) -2x -18z = -10 Inputting y = (4z -2) in equation B 2x+3y+2z=0 2x + 3*(4z-2) + 2z =0 2x + 12z -6 + 2z = 0 E) 2x +14z = 6 Adding equation D and E -4z = -4 z = 1 Now you have 1 of the 3 variables.